OpenStudy (anonymous):

A triangular pyramid has the same base and height as a triangular prism. The volume of the pyramid is 31 cubic millimeters. What is the volume of the triangular prism?

Mehek (mehek14):

prism = 3 x pyramid

OpenStudy (mathstudent55):

B is the area of the base of the prism and of the pyramid. \(\Large V_{prism} = B \times h\) \(\Large V_{pyramid} = \dfrac{1}{3} \times B \times h\) Compare the two formulas. What is the difference between them?